OpenStudy (anonymous):

Which mixed number falls between four and three-eighths and five and five-twelfths? A. four and one-fourth B. five and two-thirds C. four and one-half D. five and five-sixths

OpenStudy (mertsj):

All of those fractions can be written with a denominator of 24. That would be a good start.

OpenStudy (anonymous):

you were no help I found out the answer myself ! 4 3/8 = 4.375 and 5 5/12= 5.416 A. 4 1/4 = 4.25 B. 5 2/3 = 5.6 C. 4 1/2 = 4.5 D. 5 5/6 = 5.8 On a number line it would look like this 4.25, 4.37, 4.5, 5.4, 5.6 So the answer is C. 4 1/2

OpenStudy (mertsj):

Yw